Carbon Monoxide (CO) Mixing Ration System

 

CO      C1       CF Lamont, OK   PGS Shed at 60m Tower

 

Date

Technician

Comments

7/23/12

PAD

Mentor requests CO instrument removed and sent for repairs.  Shut down instrument at 1935 GMT.  Removed laptop, CO instrument, pump, and back pressure manifold.  All items taken to shipping dept to be sent along with (4) CO cylinders to mentor.  Laptop IBM Thinkpad, Type 0657, LBL DOE 6810776, SN LV-AV532 07/04.  CO instrument Thermo Environmental Instrument, Model 48C CO Analyzer, SN 48 CTL-78785-389, LBL DOE 6502589.  Pump KNF, Model UN026 ANI, SN 1/1041266.  External mouse and desiccant canister were also sent.  RTS at 2025 GMT. 

6/26/12

PAD

Span 1 cylinder needs to be changed for CO system.  Shut off valve at 2151 GMT and R/R cylinders.  Checked for leaks at regulator.  Old CC328854.  New CC328871.  RTS at 2151 GMT.

4/10/12

PAD

Changed CO desiccant at 1555 GMT.

3/29/12

PAD

Changed CO desiccant at 1410 GMT. 

3/22/12

CAM

Mentor notified Site Ops of abnormalities in the CO data which suggested possible water in the lines.  Checked instrument at 1845 GMT and found about 2” of water in the tub where water traps drain to.   Opened the fridge and observed quite a bit of water in the CO water trap, as well as the others, even though all the drain lines were clear and bubbling.   There was also significant moisture in the CO sample lines between the pump outlet and pressure regulator/gauge, and some visible moisture in the line going to the instrument.  Disconnected each section of tubing one at a time and let the water blow out and also blew out with canned air.  Opened up and drained all the water traps in the fridge.  RTS at 1920 GMT.  Note there has been over 2 inches of rain here this week and high humidity when not raining. Also the fridge temperature has been running around 7C with the thermostat set to max cool.  Notified mentor of findings.

2/28/12

PAD

CO span2 cylinder is down to 250 psi and needs to be replaced. Shut off cylinder and R/R at 2011 GMT.  Opened cylinder valve and checked regulator connection for leaks.  RTS at 2026 GMT.  Old CC328535, 250 psi.  New CC328541, 2,000 psi.

2/16/12

PAD

CO flow is showing 1.214 l/min.  Mentor requests flow meter replaced and IR source replaced.  Shut down system at 1916 GMT and removed power.  Opened enclosure and R/R flow meter and IR source.  Restarted system.  Flow is now showing 1.305 l/min.  AGC value is 198,319 Hz after restart for IR source.  Old flow meter, no name shown, 9934 Rev-E, SN 0842NM000089676.  New 9934 Rev-E, SN 1145NM000501326.  No information on IR source.  Troubleshooting flow issue at this time.  RTS at2046 GMT.

2/7/12

PAD

Changed CO desiccant at 1524 GMT.

1/26/12

PAD

Changed CO desiccant at 2023 GMT.

1/24/12

PAD

CO instrument flow is .918 l/min and above normal flow of .600 l/min.  Stopped labview program, rebooted computer, and powercycled instrument.  Flow returned to about 1.2 l/min after restart.  Notified mentor.  Troubleshooting at this time.  RTS at 1537 GMT.

12/13/11

PAD

R/R CO desiccant at 1852 GMT.  RTS at 1907 GMT.

12/1/11

PAD

CO span 1 cylinder is down to 290 psi and needs to be replaced.  R/R Span 1 cylinder at 2100 GMT and checked for leaks.  Old CC328952.  New CC328854.  RTS at 2115 GMT.

11/1/11

PAD

Changed CO desiccant at 1940 GMT.

9/22/11

PAD

Changed CO desiccant at 2208 GMT.

9/8/11

PAD

CO span 1 cylinder pressure is dropping out of range.  Previous troubleshooting has been unable to isolate leak.  Suspect solenoid valve inside CO analyzer.  Need to verify there are no leaks between cylinder and instrument.  Shut down instrument at 1453 GMT.  Disconnected Span 1 line from back of instrument and capped.  Cylinder currently shows 660 psi.  Will monitor to see if pressure drops with instrument disconnected.  RTS at 1507 GMT.

9/6/11

PAD

CO span 1 cylinder shows a pressure drop of 730 psi since 9/1/11.  Checked all connections at regulator and back of instrument with leak detector.  No leaks found.  Opened instrument enclosure and checked for internal leaks.  Troubleshooting at this time.  RTS at 2130 GMT.

8/12/11

PAD

Mentor requests CO pump replaced.  R/R CO pump at 2102 GMT.  RTS at 2117 GMT.  Old Model UN026 ANI, SN 1/1041267.  New Model UN026 ANI, SN 1/1041266.

8/11/11

SD

Changed CO desiccant at 2017 GMT.

8/9/11

PAD

CF lost power on 8/8/11 at 2247 GMT.  Power was restored on 8/9/11 at 0211 GMT.   CO computer is not active.  Powered up computer.  Labview program started but there is no flow on CO analyzer.  Powercycled analyzer with no change.  Rebooted computer with no change.  Shut down instrument and computer and restarted system.  Flow is now shown and instrument is operational.  Notified instrument mentor.  RTS at 1453 GMT.

8/4/11

PAD

CO span 1 cylinder is down to 200 psi and needs changed.  R/R cylinder at 1921 GMT.  Old SG9113399.  New CC328852.  RTS at 1936 GMT.

7/26/11

PAD

PGS system was removed from service.  CO instrument was shut down and relocated within the PGS shelter.  Shut CO system down at 1445 GMT and relocated after removing PGS.  RTS at 2200 GMT.

7/12/11

PAD

Changed CO desiccant at 1346 GMT.

6/14/11

PAD

Changed CO desiccant at 1822 GMT.

6/14/11

PAD

PGS vacuum pot pump has failed.  R/R vacuum pump with a spare pump for CO system.  System is now operating at 24 inHg.  RTS at 1828 GMT.

5/11/11

PAD

Changed CO desiccant at 1515 GMT.

4/7/11

PAD

R/R CO desiccant at 2104 GMT. 

3/1/11

PAD

Changed CO desiccant at 2211 GMT.

1/18/11

PAD

CO desiccant was previously changed on 1/4/11 and is changing color, being saturated with moisture, quicker than normal.  Mentor suspects a leak in the system.  R/R desiccant and found lid was not seated completely on canister.  Installed replacement canister with dry desiccant.  RTS at 1845 GMT.

1/4/11

PAD

Changed CO desiccant at 1657 GMT.

12/13/10

 

PAD

Notified by CO mentor AGC Intensity is in the low hundreds Hz.  Mentor requests infrared source changed.  Shut down instrument at 1832 GMT and R/R IR source as instructed in instrument manual.  Restarted instrument and AGC Intensity is now 200,000 Hz.  Notified instrument mentor.  RTS at 1902 GMT. 

11/18/10

PAD

AC power needed to be secured to drill pier holes for new construction near the RCF trailer.  Need to gracefully shut the instrument down prior to power outage.  Powered system down at 1420 GMT. Shut down PGS, Radon Gas, CO2 flux and CO computers. Shut down PGS IRGA rack and all pumps located in PGS enclosure. Powered down Radon Gas Analyzer unit. Restarted systems after AC power was restored. RTS at 1750 GMT.

11/15/10

PAD

CO span 1 cylinder is down to 340 psi and needs replaced.  Shut off valve and R/R cylinder at 1626 GMT.  Checked for leaks at cylinder connector and opened valve.  Old CC157152.  New SG9113399.  RTS at 1641 GMT.

11/10/10

PAD

AC power will be interrupted for construction near the main compound.  Need to gracefully shut the instrument down prior to power outage.   Powered system down at 2118 GMT.  Shut down PGS, Radon Gas, CO2 flux and CO computers.  Shut down PGS IRGA rack and all pumps located in PGS enclosure.  Powered down Radon Gas Analyzer unit.  Restarted systems after AC power was restored.  RTS at 2336 GMT.

11/9/10

PAD

Changed CO desiccant at 1549 GMT.

11/2/10

PAD

CO span 2 cylinder is down to 300 psi and needs replaced.  R/R cylinder at 2037 GMT.  Old CC20725.  New CC84203.

9/30/10

PAD

Changed CO desiccant at 1505 GMT.

8/26/10

PAD

Changed CO desiccant at 1522 GMT.

7/20/10

PAD

R/R CO desiccant at 1537 GMT.

6/22/10

PAD

Discharge tubing for CO system coming from refrigerator is blocked.  Cleared blockage with canned air at 1146 GMT.  At 1502 GMT, removed tubing between outlet connection at refrigerator and inlet connector at CO analyzer.  Blew water out of line with canned air.  Reattached connectors.  RTS at 1506 GMT.

6/11/10

PAD

Changed CO desiccant at 1443 GMT.

5/27/10

PAD

Mentor request CO pump R/R with rebuilt pump provided by mentor.  Shut off pump and R/R at 1843 GMT.  RTS at 1858 GMT.  Old SN 1/1041267.  New SN 1/1698151.

5/17/10

PAD

CF lost power on 5/16/10 between 0952-1504 GMT.  Powered on CO computer at 1250 GMT and started Labview program.  Notified mentor.  RTS at 1305 GMT.

3/25/10

PAD

R/R CO desiccant at 2023 GMT. 

3/2/10

PAD

CO system Span2 cylinder is down to 300 psi and needs replace.  At 1522 GMT, R/R cylinder.  After install, checked cylinder connector for leaks.  Old CC146248.  New CC20725. 

2/11/10

PAD

R/R CO desiccant at 1544 GMT.

12/8/09

PAD

Mentor requests CO pump swapped with spare pump supplied.  R/R CO pump at 2124 GMT.  Old SN 1/1041266.  New SN 1/1041267.  RTS at 2139 GMT.

11/10/09

PAD

Changed CO desiccant at 1635 GMT.

11/3/09

PAD

PGS computer is unable to access network.  CO and RGA computers will access.  Shut down/ restarted PGS computer.  It will now access network.  RTS at 1715 GMT.

6/16/09

PAD

R/R CO desiccant at 1614 GMT.  RTS at 1629 GMT. 

6/16/09

PAD

CO Span 2 cylinder shows 280 psi and needs replaced.  R/R cylinder at 1738 GMT.  New cylinder shows 1950 psi.   Old C29890/ New CC146248.    RTS at 1753 GMT.

5/28/09

PAD

Mentor requests CO pump swapped with rebuilt pump.  Shut down pump and R/R at 1955 GMT.  Model UN026 ANI.  Old SN 1/698151.  New SN 1/1041266.  RTS at 2010 GMT.

5/5/09

PAD

R/R CO desiccant canister at 1525 GMT.  RTS at 1540 GMT.

2/3/09

PAD

Data is suspect from CO instrument.  Mentor requests CO analyzer and laptop returned for troubleshooting and repairs.  At 1633 GMT, shut down computer and instrument and disconnected all power, comm cables, and inlet tubes.  Removed Thermo Environmental Instruments, model 48C, SN 48CTL-78785-389, LBL6502589.  Removed Toshiba laptop, model PP720U-6D90N, SN 80900062U.  Also will send second IBM Thinkpad laptop, type 0659-HQU, SN LV-AV532.

1/22/09

PAD

Mentor requests IR source replaced in CO instrument.  Shut down instrument and R/R IR source with parts and instructions provided by mentor.  Restarted instrument and computer at 1650 GMT.  Computer shows:  500 MB of free space.  Labview is unable to start until additional free space is made available.  Notified mentor who is remotely accessing computer to delete unneeded files and programs.  Unable to verify new IR source is working correctly at this time.  RTS at 1723 GMT.

1/13/09

PAD

Notified by mentor that CO data since 1/9/08 is unusable.   Mentor request instrument power cycled.  Power cycled instrument at 2032 GMT.  Mentor called later and said no improvement was made. Shut down instrument and computer and restarted.  No change in data.  Mentor is troubleshooting with vendor at this time.  RTS at 2115 GMT.

1/13/09

PAD

CO Span 1 cylinder, SN CC210306, shows 300 psi. and needs to be replaced.  Shut off line valve and removed cylinder.  Installed cylinder, SN SG9199683, checked for leaks and opened line valve.  New cylinder pressure is 1,780 psi.  RTS at 1726 GMT.

1/2/09

PAD

R/R CO desiccant at 1652 GMT.  RTS at 1707 GMT.

12/11/08

PAD

Notified by mentor CO data is unusable since power outage on 12/9/08 at 1014 GMT.  Mentor request computer rebooted and instrument power cycled.  Rebooted computer and power cycled instrument at 1940 GMT.  RTS at 1955 GMT.

10/20/08

PAD

Mentor requests CO pump swapped with replacement pump provided.  R/R CO pump at 1431 GMT.  RTS at 1446 GMT.

10/20/08

PAD

Notified via email from mentor on 10/18/08 CO instrument data is suspect.  Mentor requests computer rebooted and CO instrument power cycled.  Rebooted computer and power cycled instrument.  Notified instrument mentor.  RTS at 1427 GMT.

8/28/08

PAD

Changed CO instrument desiccant at 1946 GMT.   RTS at 2001 GMT.

8/28/08

PAD

CO drain line from is blocked.  Cleared obstruction with canned air.  RTS at 1952 GMT.

7/15/08

PAD

CF lost power on 7/13/08 between 0028 and 0137 GMT.  Notified by mentor on 7/15/08 CO computer was not collecting data.  Labview program appeared to be running correctly but would restart after trying to stop it.  Used task manager to stop program.  Rebooted computer and program restarted correctly.  RTS at 1453 GMT.

7/9/08

PAD

CF experienced power loss from 0745 to 1358 GMT.  Logged onto PGS computer.  Reset serial/USB converter to connect with datalogger.  Powered on CO2Flux computer and started Sonic1, 2, and 3 programs.  Started CO computer.  Rebooted Radon Gas Analyzer computer to reestablish comms.  Reset camera date and time.  RTS at 1515 GMT.

6/26/08

PAD

CO system Span 2 cylinder is down to 300 psi.  At 1952 GMT, shut off Span 2 cylinder, SN CC48342 / 321 ppm, and installed cylinder, SN C29890/ .324 ppm.  Checked for leaks at cylinder/ regulator connection.  RTS at 2012 GMT.

6/26/08

PAD

R/R CO instrument desiccant at 1552 GMT.  RTS at 1607 GMT.

6/10/08

PAD

Notified by mentor data files for CO system on 6/9/08 are not reasonable.  Mentor requests computer rebooted.  Rebooted CO computer at 2125 GMT.  RTS at 2140 GMT.

5/22/08

PAD

R/R CO desiccant at 1459 GMT.  RTS at 1514 GMT.

4/29/08

PAD

CF experienced extended power outage on 4/26/08 at 0034 GMT.  Power was restored at 0459 GMT.  At 1542 GMT, found CO computer has power but is not running.  Started computer and program started correctly.  RTS at 1557 GMT.

3/24/08

PAD

Notified by mentor CO system had shut down.  Computer shows error message:  Error 6 occurred at New File in Open Data File2.  Notified mentor as requested and he remotely accessed computer.  Computer hard drive was full.  Mentor deleted some data files.  Rebooted computer at 1533 and program restarted correctly.  RTS at 1543 GMT.

1/7/08

PAD

Replacement pump for CO instrument has been received at CF.  Mentor request pump installed and instrument returned to service.  At 2100 GMT removed *CO pump, according to instructions provided by mentor and installed new pump.  Powered up instrument and computer.  Opened cylinder valves for Span 1 and Span 2.  After startup, observed new pump running but instrument still displays 0 flow.  Removed inlet line and had flow at sample port from external CO pump.  Observed flow increase to .2 l/ min.  Reconnected inlet line and flow increased and stabilized at .500 l/ min.  Notified instrument mentor instrument was RTS at 2200 GMT.  *Old – 1678917   *New - 1/928230

1/2/08

PAD

At 1622 GMT, observed CO flow at 0 l/min.  External pump is running and shows 11 psi. back pressure.  Removed sample line at inlet of CO analyzer and observed flow.  Notified instrument mentor.  At 2038 GMT after consulting with mentor, removed external case of CO and observed internal pump had locked up.  Provided necessary pump information to mentor so replacement pump could be ordered.  Instrument is shut down at this time.  RTS at 2043 GMT.

10/11/07

PAD

R/R desiccant for CO instrument at 1951 GMT.  RTS at 2006 GMT.

10/3/07

PAD

CF experienced loss of power on 10/2/07.  Power has been restored and instrument needs RTS.  At 1208 GMT, turned on breakers to PGS shed.  UPS that supplies power to PGS, Radon, and CO computers shorted out.  R/R UPS with XL-1000 UPS and restored power.  Reestablished comms to PGS shed.  RTS at 1338 GMT.

3/29/07

PAD

Desiccant needs to be changed on CO analyzer.  R/R desiccant canister at 1834 GMT.  RTS at 1849 GMT.

3/16/07

PAD

Mentor request CO span 1 cylinder replaced.  At 2024 GMT, removed Span 1 cylinder CC210371, 320 psi, and installed CC210699 1990 psi.  Checked for leaks.  RTS at 2044 GMT.

3/1/07

PAD

Mentor request *CO pump changed out with spare received at CF. Shut down pump at 2027 GMT.  Removed pump, (Model UN026 ANI, SN 1/698152), and installed (Model UN026 ANI, SN 1/698151).  Restarted pump and verified all flows returned to normal.  RTS at 2047 GMT. 

7/20/06

PAD

Sebastien Biraud requests the inlet line for the NOAA sampler moved from the inlet line for the CO instrument to the 60m inlet for the PGS system.  At 1912 GMT, disconnected inlet for the NOAA and capped off the tee.  Removed cap at the 4-way cross on the 60m inlet for the PGS and attached inlet for NOAA sampler.  RTS at 1932 GMT.

5/1/06

PAD

Instrument mentor request span gas cylinder on CO instrument changed.  At 1828 GMT, removed CO span gas cylinder CC210292 that shows 300 psi, and installed cylinder CC210371.  Checked for leaks and opened all valves.  New cylinder shows 2000 psi.  RTS at 1843 GMT.

2/14/06

PAD

Mentor request *CO analyzer pump changed every 6 months.  Shut down pump, switched connectors and installed new pump.  Turned on pump at 2100 GMT.  RTS at 2115 GMT.                                              *Old - 1/698152  *New - 1/698151

12/7/05

PAD

Sebastien Biraud notified he couldn’t remotely access PGS, CO2met, CO2flux, or CO computers in PGS shed.     Reset Allied Telesyn media converters located in PGS shed and AOS trailer.  Computers can now access network.  Notified instrument mentor.  RTS at 1735 GMT.

9/6/05

PAD

 

 

 

Facilities request all computers, instruments, UPS in PGS shed be shut down so repairs can be made to conduit located between PGS shed and 60m tower.  At 1753 GMT, shut down PGS, COflux, COmet, and CO computers.  Shut down all instruments and UPS located in PGS shed.  Restarted all computers, UPS, and instruments at 1837 GMT.  PGS computer will not connect with data logger and CO2flux computer will not connect with 25m and 60m IRGA and Sonic.  Further troubleshooting indicated 25m and 60m IRGA and Sonic were operating.  Rebooted computer and power cycled port box.  PGS computer was not detecting USB hardware cable coming from data logger.  Rebooted PGS computer and it connected with data logger.  RTS at 2033 GMT. 

7/6/05

PAD

Instrument mentor notified on 7/5/05 he could not remotely access CO computer after computer was rebooted.  Windows administrator has been troubleshooting with mentor.  Tim Grove notified me that he could not remotely access computer and wanted computer checked.  Found computer had locked during Windows start up.  Power cycled computer at 1732 GMT.  Computer and Labview program restarted correctly.  Notified windows administrator.  Computer can now be remotely accessed.  RTS at 1747 GMT.

7/5/05

PAD

Sebastien Biraud notified on 7/5/05 that CO system stopped acquiring data on 6/30/05.  Found CO computer program had locked.  Computer time was correct and computer was responsive.  Rebooted computer at 2112 GMT.  Computer and Labview program restarted correctly.  Notified instrument mentor.  RTS at 2127 GMT.

6/23/05

PAD

Changed CO instrument desiccant at 1426 GMT.